a brick weighing 4 kg is dropped into a 1m deep river from a hieght of 2m . Assuming that 80% of the gravitational potential energy is finally converted into thermal energy, find this thermsl energy in calorie.

Grade:

1 Answers

Tony Abraham
9 years ago
total potential energy it had= mgh= 4*2*10 j
=80 j
its 80%=64j
its equivalent calorie=64/4.8=13.333cal.